OpenStudy (anonymous):

Try finding the slope of a line joining the following pairs of points: (x, x+2) and (x+2, x+4)

OpenStudy (anonymous):

Do you remember the equation of line slope?

OpenStudy (anonymous):

noooooooooooooooooooooooooooooooo

OpenStudy (anonymous):

It already answered in chat. But for clarity, if you know two points (x1,y1) and (x2,y2), you can find the slope of line joining that point with this formula: m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

equation is: gradient = (y2 - y1)/(x2-x1) =(x+4 - x+2)/(x+2-x) =2/2 =1